A group of eight women and ten men must select a four-person committee. How many committees are possible if more than half of the members must be women?
Answer by edjones(8007)

((10C0)*(8C4))+((10C1)*(8C3)=
=70+10*56
=70+560
=630 committees are possible.
.
Ed